State the relevant laws which are applicable on the following observation of gas: (1) Both C3H8 and CO2 have equal rate of diffusion.(2)1 dm3 O2 and 1 dm3 H2 contains same number of moles at STP.
1
Expert's answer
2017-04-22T12:34:09-0400
1) For two gases with molecular masses m1 and m2 ; and rates of diffusion R1 and R2:
R1/R2 = √m2/√m1 – (1)
2) n = v/22.4